OpenStudy (anonymous):

Which of the following is a solution to 2cos2 x – 2 = 0?

OpenStudy (anonymous):

To simplify: 2 cos 2x - 2 = 0 2 cos 2x = 2 cos 2x = 1 --- (1) To find the principal angle, do cosh 1 (cos^-1 1) Based on 1, as the value of cos is positive, find all the angles for this expression by doing the ASTC method (A and C only) A = principal angle + 2(pi)(n), where n is 0,1,2, or any other integer more than 0 C = 2pi - principal angle + 2(pi)(n), where n is 0,1,2, or any other integer more than 0

OpenStudy (anonymous):

is it 180 deegrees?

OpenStudy (anonymous):

@dauspex A-0 B-30 C-180 D-330

OpenStudy (kainui):

0=2cos(2x)–2 multiply both sides by (-1/4) 0=[1-cos(2x)]/2=sin^2(x) 0=sinx Every possible x that gives you 0 is 180n where n is any integer, so for instance for n=0, the answer is 0. For n=1, you have 180. There are literally infinite answers, but only A and C are right here.

OpenStudy (cggurumanjunath):

0 or 180